OpenStudy (anonymous):

I really need help on this i dont know what to do. Rita's company reimburses her expenses on food, lodging, and conveyance during business trips. The company pays $55 a day for food and lodging and $0.45 for each mile traveled. Rita drove 300 miles and was reimbursed $2335. Part A: Create an equation that will determine the number of days x on the trip Part B: Solve this equation justifying each step with an algebraic property of equality. Part C: How many days did Rita spend on this trip?

OpenStudy (anonymous):

55x+0.45m=reimbursement where x=days m=miles

OpenStudy (anonymous):

Is that the equation i need to solve?

OpenStudy (anonymous):

That is your equation for part A. $55*# of days 0.45c*# of miles added together gives you her payment.

OpenStudy (anonymous):

$55.45?

OpenStudy (anonymous):

No, thats just how you get the equation in part A.

OpenStudy (anonymous):

Then how do you solve that?

OpenStudy (anonymous):

55*days+0.45*300miles=2335 #days=(2335-0.45*300)/55

OpenStudy (anonymous):

i got 12733.

OpenStudy (anonymous):

im seriously no good at math im so sorry

OpenStudy (anonymous):

#days=2200/55 #days=40

OpenStudy (anonymous):

So you take $55*# of days 0.45c*# of miles as your equation then solve it and you end up getting 40 days correct?

OpenStudy (anonymous):

Yes, just substitute the values your given. Just think about it, $55 per day, so in 2 days she gets 2*55, etc. We are given the number of miles, and she gets paid $0.45 per mile, so 300*$0.45 We know she was paid $2335 total. We don't know how many days she worked for, so we get $55*days+0.45*300=2335